OpenStudy (anonymous):

Simplify 5√(7) + 12√(6) - 10√(7) - 5√(6)

OpenStudy (pottersheep):

-5√(7) + 7√(6)

OpenStudy (anonymous):

=5√7-10√7+12√6-5√6 =√7(5-10)+√6(12-5) =-5√7+7√6